Suppose you own a nice old guitar you want to sell.

To you it is valued 300. A friend of yours likes it and is willing to pay 900 euros to have it. How much is the total surplus if you sell it to your friend at 500?


1
Expert's answer
2021-01-13T10:17:10-0500
"500-300=200"

This is surplus for you? but for your friend will be


"900-500=400"

In total surplus


"200+400=600"